### v4

@@ 1 Introduction
The hard problem has sparked a substantial body of work, detailed discussion of which [14, 18] lies beyond the scope of our paper. In what follows we build upon the distinction between lower and higher order theories of information processing in relation to conscious experiences.
Higher order thought theory (HOT) [19, 20] holds that the information of which a conscious being is aware are higher order “meta-representations” of lower order “local” mental states. Lower order states may include emotions and perceptions, while higher order meta-representations reflect upon those. The link between the two may explain something of the phenomenal character of states. Sense data is processed by the body resulting in lower order mental states, and then meta representations of those is where we might find more abstract conceptual or thought-like contents of consciousness.
@@
In this paper we suggest a way to dissolve the “hard problem” of consciousness is to reverse the order and start with the ‘impure’ embodied biological organism instead of the ‘pure’ abstract mental states.
@@ 2.1 Natural Selection and Embodiment
Some have argued enactivism is incompatible with computation because the set of possible abstractions that must be searched for relevance realisation is intractable [46]. For tractability, one must first isolate a “small world” of relevant information from the “big world” of all information. Part of the reason we can address this is because we formalise embodiment, in concrete terms from the level of irreducible physical states. Embodiment implies finite resource constraints, so we adopt one more axiom from [36].
@@ 3 Relevance Realisation Through Causal Learning
Natural selection favours adaptability, hence it prefers organisms that optimise for weak policies, which we call weak policy optimisation (WPO). By constructing fit policies, an organism divides the world into relevant objects and properties. A weaker policy implies all more specific versions of itself, meaning those that more tightly constrain outputs by having a smaller extension. Hence policy learning implies a lattice of policies that vary in weakness. Every fit policy would be a “causal identity” for something relevant, for example a specific object like “this coffee” or a weaker and more general concept like “all coffee”.
@@ 4 Relevant Causal Identities
The optimal choice of policy for adaptability is the weakest because it is implied by more statements in an embodied language. That language is dispositional [61], shaped by natural selection, and so a weaker policy is likely to be a fit policy. WPO will correctly identify cause and effect where relevant, simplifying the entire environment into objects and properties relevant to the organism’s motivations.
@@ 5 Multi-Layered Self-Organisation
As the vocabulary and capacity for WPO scale, a greater variety of concepts can be learned. Progressively higher orders of ‘causal-identity’ for one’s self related information processing are constructed [34]. This lets us frame the construction of embodied selves in developmental [64] and evolutionary terms.
@@ 5.1 The First Order Self
A first order self (1ST henceforth, formalised in definition 13) allows an organism discern the consequences of its actions. This serves as the locus of self related information processing and experience [34], allowing the organism to plan complex interactions and maintain a consistent “self” that is part of the present interaction. 1ST order selves amount to reafference [65, 66], observed in mammals and insects.
@@ 6 The What and Why of Consciousness
Up to now we have developed the conceptual toolbox that we use to dissolve the hard problem. We started with the observation that human body self organises to maintain itself in the face of change both within and without. An organism does not spring into existence understanding number systems or objects, but constructs them through interaction with the environment. One can regard them as policies according to which the environment is interpreted. To learn how to interpret the environment an organism must differentiate between states, must react to change, and learn policies according to the valence associated with that change.
@@
Clearly, prelinguistic self-organising systems must classify and attach value and disvalue to states and anticipated states to prioritise and make decisions. To learn is to sense “something” and, motivated by valence, construct a policy classifying that “something”. Physical states are contentless. In the absence of language, the difference between any two states can only be qualitative. In other words we hold that quality is fundamentally valenced, i.e. the product of actively sensing and being physically attracted and repelled within a given environment.
@@
Our arguments align with those of Merker, Barron and Klein, who have linked subjective experience to reafference [72, 65, 66]. We agree that reafference is key, but we provide a very different explanation of why and how. Their work presents biological evidence for subjective experience in organisms with reafference. In contrast, we derive the 1ST order self from first principles and explain why and how it is a classifies “what it is like” to be a particular organism. The 1ST order self also happens to be equivalent to reafference, so we arrive at the same conclusion as Merker from very different, mathematical premises. Hence these are complementary positions.
@@ 8 From Rocks to Einstein: The Hierarchy of Being
To illustrate how our argument applies in the real world we describe stages of conscious organism. Each stage follows from scaling up supply and demand for WPO, through natural selection. For each stage we point out animals which are likely to be at least so conscious:
@@ 10 Appendix
(notation) $E$ with a subscript is the extension of the subscript181818e.g. $E_{l}$ is the extension of $l$..
@@
(intuitive summary) Learning is an activity undertaken by some manner of intelligent agent, and a task has been “learned” by an agent that knows a correct policy. Humans typically learn from “examples”. An example of a task is a correct output and input. A collection of examples is a child task, so “learning” is an attempt to generalise from a child to one of its parents. The lower level the child from which an agent generalises to parent, the “faster” it learns, the more sample efficient the proxy. The most sample efficient proxy is weakness [33, prop. 1, 2], which is why we’re using it here.
@@ 10.1 Proofs
Versions of proofs 1-3 were published in proceedings [33]. Further details are available in on GitHub, along with the supporting experiments [35].
@@
    From [33, prop. 3] we have proof that weakness is the optimal choice of proxy to maximise the probability of generalisation from child to parent is the weakest policy. It follows that $\mathfrak{a}$ will generalise from $\mathfrak{h}_{<t_{\mathfrak{a}}}$ to $\mu_{\mathfrak{a}}$ given the smallest history of interventions with which it is possible to do so (meaning the smallest possible ostensive definition, or cardinality $\lvert D_{\alpha}\rvert$).

### v5

@@ 1 Introduction
The hard problem has sparked a substantial body of work, detailed discussion of which [14, 18] lies beyond the scope of our paper.
In what follows we start with the observation that the mental/physical distinction that fuels the classical debates, ranging from panpsychism [19] to physicalism à la Dennett [20] may not be the only way to tackling the hard problem of consciousness. We remain agnostic to the question whether conceivability arguments such as the zombie thought experiment [21] or the knowledge argument [22, 23] are the best way to approach the problem of phenomenal consciousness. For example, one may argue that artificial systems such as Large Language Models (LLMs) conceive entities that humans never conceived or never will. It remains an open question on whether these entities exist just on the basis that a system can conceive them abstractedly or not. Our aim is more modest and starts with low level biological facts rather than high level philosophical abstractions to reformulate the problem of consciousness in a way that does not take for granted the mental/physical division in the first place. In this paper we suggest a way to dissolve the “hard problem” of consciousness is to reverse the order and start with the ‘impure’ embodied biological organism instead of the ‘pure’ abstract mental states.
Take for example the higher order thought theory (HOT) [24, 25] holds that the information of which a conscious being is aware are higher order “meta-representations” of lower order “local” mental states. Lower order states may include emotions and perceptions, while higher order meta-representations reflect upon those. The link between the two may explain something of the phenomenal character of states. Sense data is processed by the body resulting in lower order mental states, and then meta representations of those is where we might find more abstract conceptual or thought-like contents of consciousness.
@@ 2.1 Natural Selection and Embodiment
Some have argued enactivism is incompatible with computation because the set of possible abstractions that must be searched for relevance realisation is intractable [51]. For tractability, one must first isolate a “small world” of relevant information from the “big world” of all information [52]. Part of the reason we can address this is because we formalise embodiment, in concrete terms from the level of irreducible physical states. Embodiment implies finite resource constraints, so we adopt one more axiom from [41].
@@ 3 Relevance Realisation Through Causal Learning
Natural selection favours adaptability, hence it prefers organisms that optimise for weak policies, which we call weak policy optimisation (w-maxing). By constructing fit policies, an organism divides the world into relevant objects and properties. A weaker policy implies all more specific versions of itself, meaning those that more tightly constrain outputs by having a smaller extension. Hence policy learning implies a lattice of policies that vary in weakness. Every fit policy would be a “causal identity” for something relevant, for example a specific object like “this coffee” or a weaker and more general concept like “all coffee”.
@@ 4 Relevant Causal Identities
The optimal choice of policy for adaptability is the weakest because it is implied by more statements in an embodied language. That language is dispositional [67], shaped by natural selection, and so a weaker policy is likely to be a fit policy. w-maxing will correctly identify cause and effect where relevant, simplifying the entire environment into objects and properties relevant to the organism’s motivations.
@@ 5 Multi-Layered Self-Organisation
As the vocabulary and capacity for w-maxing scale, a greater variety of concepts can be learned. Progressively higher orders of ‘causal-identity’ for one’s self related information processing are constructed [39]. This lets us frame the construction of embodied selves in developmental [70] and evolutionary terms.
@@ 5.1 The First Order Self
A first order self (1ST henceforth, formalised in definition 13) allows an organism to discern the consequences of its actions. This serves as the locus of self related information processing and experience [39], allowing the organism to plan complex interactions and maintain a consistent “self” that is part of the present interaction. 1ST order selves amount to reafference [71, 72], observed in mammals and insects.
@@ 6 The What and Why of Consciousness
Up to now we have developed the conceptual toolbox that we use to dissolve the hard problem. We started with the observation that a human body self organises to maintain itself in the face of change both within and without. An organism does not spring into existence understanding number systems or objects, but constructs them through interaction with the environment. One can regard them as policies according to which the environment is interpreted. To learn how to interpret the environment an organism must differentiate between states, must react to change, and learn policies according to the valence associated with that change.
@@
For biological, living self-organising systems, being alive is intrinsically good, being dead or ill is bad. For example, in developing self-organising systems such as the human body, it is good to have a better developed sense of smell than vision at early stages of life (babies are more accurate smellers than adults). Different senses have different degrees of valences at different stages of the developmental ladder. Hence there is “something what it is like” to be a baby which is different from something what it is like to be an adult. However, the idea here is that basic experiences (i.e. the self-organising process of staying alive) have valence at their very core (good=stay alive; bad=dead). This culminates in experiential, phenomenal ‘quality’ of appearance, of what it is likeness and how thing appear to the system (i.e. phenomena). That is the aspect that philosophers typically tackle. Our aim is not to dismiss the latter; our claim is that one cannot have the latter without the former.
Clearly, prelinguistic self-organising systems must classify and attach value and disvalue to states and anticipated states to prioritise and make decisions. To learn is to sense “something” and, motivated by valence, construct a policy classifying that “something”. When we say physical states are contentless, we are not suggesting physical objects and properties don’t exist. Objectively, a physical object is an aspect of the environment and every aspect is realized by one state or another. However, previous work has shown that for every aspect of the environment there is an equivalent program, meaning from an objective point of view what is or is not a physical object or property is a matter of interpretation [41]. Subjectively an object, property or quality only exists if an organism constructs a causal identity for it. At the most basic level, the foundation of all of this is physical attraction and repulsion. An organism hard-wired by natural selection can be physically attracted or repelled without any concept of what it is physically attracted or repelled from. It doesn’t know to what it is attracted or repelled from in the sense that it has not constructed a causal identity for that object. There is no named object which has the property ‘attractive’. The organism itself is simply attracted or repelled according to hard-wired policy. For example, a single-celled organism might be attracted to glucose, and its response to ‘tumble’ or ‘stop’ would be hard-wired. It does not satisfy the incentive or scale preconditions to construct a causal identity for glucose, or ‘sweet’, or hunger, or anything else. Intuitively, one might think of this like ‘one dimensional’ valence. To the bacteria, there is no subjective categorical variable or object to which it is attracted, though in the objective sense such a thing does exist. It is simply attracted or repelled. Over the course of this paper we have sought to explain how we get from simple ‘one dimensional’ hard-wired valence to a property or quality that has valence. Categorical variables for properties or ‘qualities’ like ‘the colour red’ or ‘the smell of coffee’ only exist when we have a causal identity for them, which requires scale and incentive. It can be hard-wired, or learned. When we have only one cell we do not have a vocabulary capable of satisfying the scale precondition. However self-organising biological systems are collective [78]. Control is delegated and distributed among smaller components [43] which operate concurrently. An individual cell might only have ‘one dimensional’ valence, but as soon as we have two cells we have something analogous to a second ‘dimension’. Cellular collectives are polycomputational systems, meaning one cell can play a part in more than one computation simultanously, at different scales [79] [^15] This is what the weakness of a policy captures; the number of interactions of which it can be a part.. As we scale up the collective, its is impelled by the cells of which it is formed. Instead of ‘one dimensional’ valence, we now have a rich tapestry of competing drives. It is this rich tapestry that is a property or ‘quality’ of a state of the environment interpreted by the organism. It is this rich tapestry that can be reduced to causal identities for properties like ‘the color red’, ‘hunger’ or ‘thirst’. A higher level of abstraction is formed from this rich tapestry of valence and how it changes over many interactions (the different inputs and outputs in a task that imply generalizable causal identities). Hunger and thirst might have the same overall intensity and so one could say they have the same overall valence, but they are qualitatively different because they are different ‘tapestries’ of valence at a lower level of abstraction. This is what is meant by The Psychophysical Principle of Causality in section 3. In a self-organising system of sufficient scale with sufficient incentive, what starts out as simple ‘one dimensional’ valence culminates in categorical variables that serve different organismic needs. This is why the sound of middle C is different from the color blue. They have different causal identities. Middle C results from sound waves hitting the ear, blue results from light waves hitting the eye. These variations produce different qualia. Qualia do not need to be perfect. They simply need to be “good enough” to motivate actions necessary for survival in the environment of the organism. Qualia are produced by the sensory and nervous systems of the body. The valence of red light waves misses the eyes of dogs entirely. Dogs rely more on smell than sight. This suits their environmental niche. At a sufficient scale causal identities can be constructed, stored and recalled by their quality and valence.
@@
Our arguments align with those of Merker, who has linked subjective experience to reafference [80, 71, 72]. We agree that reafference is key, but we provide a very different explanation of why and how. Their work presents biological evidence for subjective experience in organisms with reafference. In contrast, we derive the 1ST order self from first principles and explain why and how it is a classifies “what it is like” to be a particular organism. The 1ST order self also happens to be equivalent to reafference, so we arrive at the same conclusion as Merker from very different, mathematical premises. Hence these are complementary positions.
@@ 8 From Rocks to Einstein: The Hierarchy of Being
To illustrate how our argument applies in the real world we describe stages of conscious organism. Each stage follows from scaling up supply and demand for w-maxing, through natural selection. For each stage we point out animals which are likely to be at least so conscious:
@@ 10 Appendix
(notation) $E$ with a subscript is the extension of the subscript191919e.g. $E_{l}$ is the extension of $l$..
@@
(intuitive summary) Learning is an activity undertaken by some manner of intelligent agent, and a task has been “learned” by an agent that knows a correct policy. Humans typically learn from “examples”. An example of a task is a correct output and input. A collection of examples is a child task, so “learning” is an attempt to generalise from a child to one of its parents. The lower level the child from which an agent generalises to parent, the “faster” it learns, the more sample efficient the proxy. The most sample efficient proxy is weakness [38, prop. 1, 2], which is why we’re using it here.
@@ 10.1 Proofs
Versions of proofs 1-3 were published in [38]. Further details are available in on GitHub, along with the supporting experiments [40].
@@
    From [38] we have proof that weakness is the optimal choice of proxy to maximise the probability of generalisation from child to parent is the weakest policy. It follows that $\mathfrak{a}$ will generalise from $\mathfrak{h}_{<t_{\mathfrak{a}}}$ to $\mu_{\mathfrak{a}}$ given the smallest history of interventions with which it is possible to do so (meaning the smallest possible ostensive definition, or cardinality $\lvert D_{\alpha}\rvert$).